OpenStudy (anonymous):

simplify the given expression. (8x/4y) x (17y^2/16x^3)

OpenStudy (anonymous):

\[(\frac{8x}{4y})\times(\frac{17y^{2}}{16x^{3}})=\frac{8x^{1-3}\times17y^{2-1}}{4\times16}\]